OpenStudy (anonymous):

Factorize these Quadratic Expressions; 1) 3x^2 - 12, 2) 10w^2 - 1000, 3) 3^2-100, 4) x^2 - x - 12, 5) x^2 - 3x - 18.

OpenStudy (jamesj):

I'll do the first one for you and you can then figure out how to do 2 and 3. 3x^2 - 12 = 3x^2 - 3(4) = 3 (x^2 - 4) Now x^2 - 4 = x^2 - 2^2 a difference of squares, hence x^2 - 4 = (x-2)(x+2) Thus 3x^2 - 12 = 3(x-2)(x+2)

OpenStudy (anonymous):

(x-6)(x+3) is the last one

OpenStudy (anonymous):

(x-4)(x-3) is the 4th one

OpenStudy (anonymous):

10(w-10)^2 is the second one and im not sure if you wrote the 3rd one wrong, if not its -91